Pass a journal entries in the in recognition of both an asset and a liability in the books of the lessee at the inception of the lease at amount equal to present value of minimum lease payments.
Answer #1

IFRS 16 Leases brings significant changes in accounting requirements for lease accounting, primarily for lessees.

IFRS 16 replaces the existing suite of standards and interpretations on leases:

– IAS 17 Leases

– IFRIC 4 Determining whether an Arrangement contains a Lease

– SIC 15 Operating Leases – Incentives

– SIC 27 Evaluating the Substance of Transactions Involving the Legal Form of a Lease

Lease Liability – Initial Recognition

As per IFRS 16

Lease Liability/Minimum Lease Payments = Fixed Payments from commencement date + Certain Variable Payments + Residual Value Guarantee +  Exercise Price of Purchase Options + Termination penalties

Right of Use Asset = Lease Liability + Initial Direct Costs + Costs of removal and restoring + Payments made at or prior to commencement Lease + incentives received

Journal Entries with an Basic Example.

Lessee enters into a 5-year lease of a Equipment .Economic Useful Life of asset is 5 Year.Annual Lease payments of $30,000 at end of year and Interest Rate @ 5% .Lease Covey No Ownership at the end of Period ,no purchase option at the end of Lease & no Residual Guarantee Value. Because Lease is for full life of and Asset .Lease will be considered as Finance Lease as IFRS 16. Depreciation at Straight Line Method.

Solution :

Present Value of Annuity for 5 Year @ 5% = 4.32948

Present Value of Lease Payment will be as Under : $30000 * 4.32948 = $129,884

Year Date Lease Paymnent Interest @5% Interest Amount Amortization of Lease Liablity Finance Lease Liablity
1 1St Jan - -                               1,29,884
1 31st Dec $30,000 129884*5% = 6,494                                                       23,506                               1,06,378
2 31st Dec $30,000 106378*5% =                                                    5,319                                                       24,681                                   81,697
3 31st Dec $30,000 81697*5% =                                                    4,085                                                       25,915                                   55,782
4 31st Dec $30,000 55782*5% =                                                    2,789                                                       27,211                                   28,571
5 31st Dec $30,000 28571*5% =                                                    1,429                                                       28,571 -

Journal Entries Will be as UNDER :

Year Date Dr/Cr Journal Entries Amount Remark
1 1st Jan Debit Finance Lease - Right-of-use Asset                                              1,29,884
Credit Finance Lease Liablity                                              1,29,884
1 31st dec Debit Interest Expense                                                    6,494
Debit Finance Lease Liablity                                                 23,506
Credit Cash/Bank                                                 30,000
1 31st dec Debit Depreciation Expense                                                 25,976 Straight Line - $129884/5 year
Credit Accumulated Depreciation - Finance Lease - Right-of-use Asset                                                 25,976